# Coupling induced emergent topology in a two-leg fermionic ladder

### v1

@@ V Effect of uniform nearest-neighbor interaction
Figure 15: The figure shows the magnitude of bond energies of all the bonds defined in Eq.(13) and the on-site particle density for different phases corresponding to Fig. 9 with a system consisting of $L=100$ rungs. Only the central $10$ rungs of a system with $L=100$ rungs are shown to focus on the bulk behavior. In the figures, the thickness of a bond is proportional to the respective magnitude of bond energy, and the face color of the circles represents the values of the onsite particle density for (a) $V=0.2$, and (b) $V=2.8$.
@@
$\hat{H}{{}_{U}}=\hat{H}+\hat{H}_{V_{U}}$ (17)
where $\hat{H}$ is the non-interacting part described earlier, and the interaction term $\hat{H}_{{V_{U}}}$ is defined as
$\hat{H}_{{V_{U}}}$ $=V\sum_{i}\left(\hat{n}_{i,a}-\frac{1}{2}\right)\left(\hat{n}_{i,b}-\frac{1}{2}\right)$
@@
VI Conclusion
In this study, we explored the effect of inter-leg coupling in establishing topological phases of spinless fermions on a two-leg ladder. By assuming one of the legs dimerized in nearest-neighbour hopping and the other uniform, we showed that the system as a whole becomes topological as soon as the inter-leg hopping is switched on. However, with increase in the inter-leg hopping, the system undergoes a transition to a trivial phase. We also found that this emergent topological phase in the ladder is independent of the dimerization pattern in the dimerized hopping leg. As a result, we obtained two different topological phases of opposite character. Crucially, we showed that the inter-leg interactions can induce a distinct topological phase transition, revealing an interaction-induced topological phase transition in the system.
We characterized this transition by multiple diagnostics, including the entanglement spectrum, topological invariants, string order parameters and Thouless charge pumping. Our study reveals a non-trivial topological phenomenon involving spinless fermions on a quasi-1D geometry which clearly highlights the intricate interplay between lattice geometry, many-body effects and particle statistics in stabilizing the topological phase transitions in coupled systems. This also opens up possible avenues to explore the fate of these topological phases in bosonic and spinful fermionic many-body systems, in lattices other than ladder geometries, and in the presence of disorder.
VII Acknowledgment

### v2

@@ III.2.2 Effect of uniform interaction
$\hat{H}{{}_{U}}=\hat{H}+\hat{H}_{V}$ (17)
where $\hat{H}$ is the non-interacting part described earlier (Eq. (2)), and the interaction term $\hat{H}_{{V}}$ is defined as
$\hat{H}_{V}$ $=V\sum_{i}\left(\hat{n}_{i,a}-\frac{1}{2}\right)\left(\hat{n}_{i,b}-\frac{1}{2}\right)$
@@
Figure 15: The figure shows the magnitude of bond energies of all the bonds defined in Eq. (13) and the on-site particle density for different phases corresponding to Fig. 14 with a system consisting of $L=100$ rungs keeping $t_{1}=0.2,~t_{2}=1.0,~t_{p}=0.8$ fixed. Only the central $10$ rungs of a system with $L=100$ rungs are shown to focus on the bulk behavior. In the figures, the thickness of a bond is proportional to the respective magnitude of bond energy, and the face color of the circles represents the values of $\langle\hat{n}_{i,\alpha}\rangle$ for (a) $V=0.2$, and (b) $V=2.8$.
IV Conclusion
In this study, we have explored the effect of inter-leg coupling in establishing topological phases and phase transitions of spinless fermions on a two-leg ladder. By assuming that one of the legs possessing nearest-neighbour hopping dimerization and the other leg a uniform chain, we have shown that the entire ladder exhibits topological character irrespective of the dimerization pattern in the dimerized leg. As a result of this, we obtained two different topological phases of opposite character by changing the dimerization pattern of the leg possessing dimerized hopping. However, stronger inter-leg hopping makes the entire system a trivial insulator. Additionally, we showed that the inter-leg interactions can induce a distinct topological phase transition, revealing an interaction-induced topological phase transition in the system. We characterized these transitions by multiple diagnostics, including the entanglement spectrum, topological invariants, string order parameters and Thouless charge pumping.
Our study reveals topological phenomena involving spinless fermions on a quasi-1D geometry, which clearly highlights the intricate interplay between lattice geometry, many-body effects, and particle statistics in stabilizing the topological phase transitions in coupled systems. This also opens up possible avenues to explore the fate of these topological phases in bosonic and spinful fermionic many-body systems, in lattices other than ladder geometries, and in the presence of disorder.
V Acknowledgment
@@
VI Data availability
The data that support the findings of this article are openly available Parida et al. .